5 out of 15 or 1/3 1:3
What is the chance of drawing a 6 form a deck of cards?
There are four 6s in a standard deck of 52 cards, so the chance of drawing a 6 is:
4/52 = 1/13 = 0.0769
Suppose you roll 2 dice. what is the probability that you will roll a 3 or 5?
The probability that you roll a 3 or a 5 on one or both dice is 20/36 = 5/9
What would the theoretical probability be of flipping a coin four times?
Anyone can flip a coin four times so I say 100 percent probability. On the other maybe you should ask the odds of the results from four flips.
Prob(At least one ace in ten cards out of 52)
= 1 - Prob(No aces in ten cards out of 52)
When you start there are 48 non-ace cards out of 52,
At the next draw, there are 47 non-ace cards out of 51, and so on
= 1 - (48/52)*(47/51)* ... *(39/43) = 1 - 0.41 = 0.59
What is the probability of obtaining tails and a five?
The question is not specific enough for an answer.
If you do not provide the necessary information, asking such a question is pointless.
What are you looking for when you are finding the mode of a set of data?
The mode is the data point that occurs the most number of times; in addition the data could be bimodal (2 modes) or multimodal (3 or more modes).
What is the mode in the data sequance 30 10 40 20?
The mode in a set of data is the number which appears most frequently. For instance in the data set 2,2,3,5,6,7,7,7,7,8,9,9, the mode is clearly 7 as it appears most frequently. If there are two numbers that appear equally frequently, such as in the set 1,2,2,3,4,4,4,4,5,6,6,7,8,9,9,9,9, we would say the data is bimodal, and has modes of 4 and 9. If there are 3 equally frequent values, then the data is trimodal and so on. In your example all numbers are equally frequent (each occurring once). In this case we would say there is no mode.
What are the disadvantages of probability sampling?
The three main disadvantages are non-response, representation and expense.
Under probability sampling once the sample is selected the researcher may not interfere with it sample. Specifically, if a large proportion of the sample do not respond, the study will be seriously flawed. With quota sampling, for example, the sampling will continue until the required number of responses are obtained.
The representation issue arises if a small sample is taken from a large population which is made of of sub-populations. Elements within each sub-population are similar to one another but there are important differences between members of different sub-populations. Some sub-populations may end up not being represented in the sample. A typical school-level example is sampling all employees of a company: 2 managers, 90 workers, 8 office staff. If using 10% samples, most (96% of them) will exclude any managers. Stratified sampling will get around that.
The expenses issue can be best illustrated by an example. Suppose you want to sample 1% of all schools in your country. You get a list of all schools in the country and select a 1% random sample. These will be scattered all over the country and you (or your team) will spend a huge amount of time and money travelling from one location to another. A better alternative is to divide up the country into a number of regions, each with [approximately] the same number of schools. Select a 1% sample of the regions [not schools] and then conduct a census of all schools within the selected regions. A variant might be to sample 2% of the regions and, within each region sample 50% of the schools so as to give a 2%*50% = 1% sample overall.
What does unfavourable outcome mean?
It is an outcome of a trial in which the event of interest does not occur.
What is the probability of drawing two cards from a deck and they both are queens?
It is approx 0.009050
What is the event space for rolling the sum of 8 with two dice?
There are 36 permutations of rolling two dice. Of these, there are five that add up to a sum of 8...
6+2
5+3
4+4
3+5
2+6
This translates to an event (or sample) space of {62, 53, 44, 35, 26}.
The probability, then, of rolling a sum of 8 on two dice is 5 in 36, or about 0.1389.
What are the odds of getting a full house in 7 card poker if your first 3 cards are queens?
If our first three cards are queens, our question really is, what is the probability that i get a pair given four cards out of 49? Since we already have three queens, we don't want to consider the case where we can get a pair of queens, being as this is clearly impossible. There are four of each card in every suit, so the number of ways we can get any pair is 4C2*12*4C2, or 432. The total number of ways we can get any two cards is 49C2, or 1176. Therefore, our probability that we get a pair in any two places of those four cards is 432/1176, or 18/49. However, we aren't done yet because we have to make sure that we don't get another queen or two more of the same previous two cards that we drew on the remaining two cards, because either case would result in four of a kind. Therefore, we need to multiply 18/49 by the odds that neither of the two previous instances occur. The probability that they don't occur is 1-(probability that they do occur). So, the number of ways to get the queen is 2C1*1 or 2, and the number of ways to draw the same two cards is 2C2, or 1. Our total number of ways to draw 2 cards from 47 is 47C2, or 1081. So, the odds that we get either the other queen or the other two cards is 3/1081. To get the probability that our cards do not result in either of those two instances is 1-3/1081, or 1078/1081. So, our final probability is 18/49*1078/1081, or 396/1081, giving you an approximate 36.6 percent chance of getting a full house.
For decades the go to place to find the best quality reviews on pretty much any product in the US has been consumerreports.org. Consumer reports has been around for forever, and they review products that come onto the market and give pretty solid objective reviews. They can help make your decision easier.
There is a probability of 3 that it will be blue.
1000 Hz. The others are outside the range of human hearing. You can feel 10 Hz but not hear it. C. 1,000 Hz
What is the formula to calculate size of a sample?
by using the capture-recapture method which gives you the total size of organisms in a population.
What is the Absolute number of Powerball variations?
Powerball is a game of combinations of 59 things (balls) taken 5 at a time, plus the combinations of 39 things taken 1 at a time.
The combinations of N things taken P at a time is (N! - P!) / P!
For the first part, this is 5,006,386. For the second part, this is 39. The total number of combinations is the product of the two, or 195,249,054.
The Australian version is similar, but uses 45 things in both parts, so the total combinations is 1,221,759 times 45, or 54,979,155.
In a standard deck of 52 playing cards how many cards are a five or a nine?
There are four fives and four nines; so the number of cards that are a five or a nine is 4 + 4 or 8.
What is the probability the number contains at least one digit 2?
The probability is 1. And the same is true for any other digit.
This counter-intuitive result can be proved as follows:
Ignore numbers with leading zeros.
There are 9 1-digit numbers of which 8 do not contain any 2s;
There are 90 2-digit numbers of which 8*9 do not contain any 2s;
There are 900 3-digit numbers of which 8*9*9 do not contain any 2s;
and so on.
The total number of k digit numbers is 9 + 90 + 900 + ... + 9*10^(k-1)
= 9*[10^k-1]/[10 - 1] = 10^k - 1
Of these, the count of numbers which do not contain any 2s = 8 + 8*9 + 8*9^2 + ... + 8*9^(k-1)
= 8*[9^k-1]/[9 - 1] = 9^k - 1
So the probability that the numbers do not contain any 2s is [9^k - 1]/[10^k - 1] which, as k gets larger, approaches 9^k/10^k = (0.9)^k which tends to 0.
So, the probability that the number DOES contain at least one 2 = 1 - Pr(It contains none) = 1 - 0 = 1.
What percents of cards are diamonds in a standard deck of cards?
Not counting jokers, there are equal amounts of diamonds, hearts, spades, and clubs. 1/4 of the cards are diamonds- 25%.
0.05 I think is the answer
What is the probability of getting one tail if you tossed 3 coins?
Three coins can come up in (2 x 2 x 2) = 8 different ways.
One tail can come up in 3 different ways . . . one for each coin.
The probability of one tail is (3/8) = 0.375 = 37.5%